# Modeling local predictive ability using power-transformed Gaussian processes

### v1

@@ 4.1 Inferring local predictive ability
The bike sharing dataset in Fanaee-T and Gama, (2014) includes the daily number of rentals, our main variable of interest, from January 1, 2011 to December 31, 2012. To predict this variable, the experts use several covariates related to the weather, an indicator for season, and the number of bike rentals the previous day. They also use indicators for workday and holiday, the latter being based on a list of official US holidays.
@@
We use the same setup as in Oelrich et al., (2021), with three experts: a Bayesian regression model (BREG), a Bayesian additive regression tree model (BART), and a Bayesian linear regression model with stochastic volatility (SVBVAR). For further discussion of the pooling variables and the data set, see Oelrich et al., (2021).
@@
To estimate local predictive ability, we use the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ method, as the number of data points makes the $\operatorname{GP}(\chi^{2}_{1})$ method inconveniently slow. More specifically, we use a GP with a squared exponential kernel with automatic relevance determination (ARD) so that each pooling variable is associated with its own length scale hyperparameter (Rasmussen and Williams,, 2006). Following the Gaussian process regression example in the Stan User’s Guide, we give the length scales independent $\operatorname{Inverse-Gamma}(5,5)$ priors, and the noise and signal variance are given $\mathcal{N}^{+}(0,1)$ priors (Stan Development Team,, 2022).
@@
Figure 7: Estimates of local predictive ability for a subset of the bike sharing data. The black line shows the observed log scores and the red ribbon shows a $95\%$ HPD interval for the posterior predictive distribution of log scores. See Appendix A for plots of the full data set.
@@
Figure 7 shows a $95\%$ highest posterior density (HPD) interval for the predictive distribution of a subset of the log scores; see Appendix A for results for the full dataset. For all experts, the majority of observed log scores land within the HPD. However, around twice as many as would be expected fall outside, with BART doing the worst and SVBVAR the best. This may indicate that the decision maker could be missing an important pooling variable.
@@
 global meansuperscript $\ell^{\prime}\text{ global mean}$ $:\hskip 14.22636pt\ell^{\prime}_{ik}\sim\mathcal{N}(\mu,\sigma^{2}),$ (19)
@@
4.2 Using local predictive ability for forecast combination
Inferred local predictive ability has a natural use when combining expert forecasts. A natural approach is to weigh each expert directly by the posterior probability that the predictive ability of that expert is greater than all other experts, $p\big{(}\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}})$). In practice however, when the amount of data is small these weights will tend to not discriminate strongly enough between experts and be outperformed by more aggressive weighing schemes, which leads us to consider other methods. As a flexible aggregation weights based on local predictive ability we propose feeding $p\big{(}\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}})$) through a softmax transformation
$w_{k}(\tilde{\mathbf{z}})=\frac{\exp\Big{(}c\times p\big{(}\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}})\big{)}\Big{)}}{\sum_{k=1}^{K}\exp\Big{(}c\times p\Big{(}\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}})\Big{)}\Big{)}.}$ (20)
@@ 4.2 Using local predictive ability for forecast combination
When the decision maker has no strong prior opinion of $c$, it can be dynamically selected at each time point by selecting the value that optimizes the historical predictive ability of the pool. We refer to this method as $\operatorname{GP}(\nicefrac{{1}}{{3}})$ dynamic in tables and figures.
Two alternatives to this method that we will use as benchmarks is to simply weigh by the posterior probability $p(\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}}))$, or to assign weight one to the model with the highest posterior expectation of $\eta(\tilde{\mathbf{z}})$. We will refer to these two variants in tables and figures as $\operatorname{GP}(\nicefrac{{1}}{{3}})$ natural and $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model selection, respectively.
@@
Figure 10 shows how the optimal discrimination changes over time for the bike share data. At each time point, the historical performance of the pool for $0\leq c\leq 20$ is evaluated, and the next prediction is made with the value of $c$ corresponding to the historically most accurate pool. The optimal discrimination $c$ is volatile when little historical data is available, but then settles down around $c=5$ as a compromise between the extremes of equal weights and assigning all weight to the expert with highest probability of having the best predictive ability at $\tilde{\mathbf{z}}$.
Figure 11 shows the cumulative log predictive density for the three methods. The dynamic method perform best, closely followed by the selection method.
Figure 10: Optimal value of the discrimation factor $c$ in (20) over time in the bike rental data.
Figure 11: Cumulative log predictive density for three methods based on local predictive ability. The legend is sorted in descending order of total log predictive score.
@@
Figure 12 compares the best performing of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ methods with a selection of reference methods: the caliper method used in Oelrich et al., (2021), which is a nonparametric method that averages log scores of nearby observations in $\mathbf{z}$-space; the global optimal pool of Geweke and Amisano, (2011) which selects aggregation weights that optimize the historical performance of the pool in terms of log scores; the local optimal pool in Oelrich et al., (2021), which is an extension of the optimal pool of Geweke and Amisano, (2011) where the optimal weights are determined on a subset of the data in the pooling space closest to the current point. For more detail about these reference methods, see Oelrich et al., (2021). Figure 12 shows that the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ initially struggles when there is little training data, but then quickly improves over time and at the end of the sample it only beaten by the local optimal pool. The total sums for all methods, together with the performance of the individual experts, are shown in Table 2.
@@
Method Sum of log scores
SVBVAR 45.06
BREG 28.76
BART 43.64
Equal weights 98.65
Caliper 133.66
Global opt. 130.15
Local opt. 145.27
$\operatorname{GP}(\nicefrac{{1}}{{3}})$ natural 107.12
$\operatorname{GP}(\nicefrac{{1}}{{3}})$ model sel. 126.62
$\operatorname{GP}(\nicefrac{{1}}{{3}})$ dynamic 136.44
@@ 5 Conclusions and further research
Gaussian processes have the drawback of being computationally heavy, and so we suggest the use of a power transformation of log scores which allows us to use a much faster marginalized GP to generate draws from the posterior of $\eta(\mathbf{z})$. Our proposed method is shown to work well for predicting bike share data (Fanaee-T and Gama,, 2014), where it outperforms all compared methods except the local version of the linear prediction pool of Geweke and Amisano, (2011) in Oelrich et al., (2021).
Our method models the marginal distribution of the predictive ability $\eta_{k}(\mathbf{z})$ for each expert. Extending the method to model the predictive ability of the experts jointly, for example using a multi-output GP, is a natural extension, and is likely to have an impact on the pooling of experts (Winkler,, 1981). Further work should also explore the properties of the general power transformed Gaussian process model proposed here, as well as deriving the distribution of log scores under different assumptions of expert distribution and data generating processes.
@@
Our paper focuses on the effect of pooling variables and therefore deliberately ignores any persistence of local predictive ability over time. Allowing for time-varying weights is shown in Del Negro et al., (2014); Billio et al., (2013); McAlinn et al., (2020) to be beneficial in macroeconomic forecasting. While it is straightforward to include some function of time in the vector of pooling variables in our framework, some care should be taken to define an appropriate kernel function with respect to time.
Appendix A Estimates of local predictive ability
Figure 13: Estimates of local predictive ability for the BART model. The black line shows the observed log scores and the red ribbon shows a $95\%$ HPD interval for the posterior predictive distribution of log scores.
Figure 14: Estimates of local predictive ability for the BREG model. The black line shows the observed log scores and the red ribbon shows a $95\%$ HPD interval for the posterior predictive distribution of log scores.
Figure 15: Estimates of local predictive ability for the SVBVAR model. The black line shows the observed log scores and the red ribbon shows a $95\%$ HPD interval for the posterior predictive distribution of log scores.

### v2

@@ 3.1 Data and models
The dataset in Fanaee-T and Gama, (2014) includes the daily number of bike rentals over the two year period from January 1, 2011, to December 31, 2012. The dataset further contains several covariates relating to the weather, as well as indicators for official US holidays and an indicator for if the day in question is a workday.
@@
We use the same three experts as in Oelrich et al., (2024) to generate predictions: a Bayesian regression model (BREG), a Bayesian additive regression tree model (BART), and a Bayesian linear regression model with stochastic volatility (SVBVAR). These experts use several weather-related variables, an indicator for season, the number of rentals the previous day, and indicators for workday and holiday as covariates. For further discussion of the data see Oelrich et al., (2024).
As pooling variables the decision maker uses humidity, wind speed, and temperature from the dataset in Fanaee-T and Gama, (2014). The decision maker also decides to use a variable she calls family holiday which is not included in the dataset, or in the estimation of any of the experts, which takes the value $1$ at Thanksgiving and Christmas (Eve and Day) (Oelrich et al.,, 2024). This variable is included to represent the decision makers belief that there are certain stay-at-home holidays where the demand for rental bikes is almost non-existent. As neither of the experts have access to this variable, she expects their predictive ability to vary significantly with regards to this variable.
3.2 Estimating predictive ability
@@
To estimate local predictive ability, we use the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ method, as the number of data points makes the $\operatorname{GP}(\chi^{2}_{1})$ method inconveniently slow. More specifically, we use a GP with a squared exponential kernel with automatic relevance determination (ARD) so that each pooling variable is associated with its own length scale hyperparameter (Rasmussen and Williams,, 2006). Following the Gaussian process regression example in the Stan User’s Guide, we give the length scales independent $\operatorname{Inverse-Gamma}(5,5)$ priors, and the noise and signal variance are given $\mathcal{N}^{+}(0,1)$ priors (Stan Development Team,, 2024).
@@
Figure 6 shows a $95\%$ highest posterior density (HPD) interval for the predictive distribution of a subset of the log scores. For all experts, the majority of observed log scores land within the HPD. However, around twice as many as would be expected fall outside, with BART doing the worst and SVBVAR the best. This may indicate that the decision maker could be missing an important pooling variable.
Figure 6: Estimates of local predictive ability for a subset of the bike sharing data. The black line shows the observed log scores and the red ribbon shows a $95\%$ HPD interval for the posterior predictive distribution of log scores.
@@
′′ global meansuperscript′′ $\ell^{\prime\prime}\text{ global mean}$ $:\hskip 14.22636pt\ell^{\prime\prime}_{ik}\sim\mathcal{N}(\mu,\sigma^{2}),$ (25)
@@
Figure 8: Joint predictive distributions of log scores for the last data point in the bike-sharing data. The graphs show kernel density estimates for univariate (left side) and multivariate (right side) Gaussian processes. The observed log scores are marked with a red dot.
To estimate the joint predictive ability of all three experts we use the $\operatorname{multi-GP}(\nicefrac{{1}}{{3}})$ model, also with a squared exponential automatic relevance determination (ARD) kernel (Rasmussen and Williams,, 2006). To allow for the possibility that some pooling variables are unimportant to some experts we use a fairly flat prior for the length scales, $\operatorname{Cauchy}(0,5)$, restricted to the interval $(0,100)$. For the $\mathbf{C}$ matrix we use a $\mathrm{LKJ}(\eta=3)$ prior (Lewandowski et al.,, 2009), and zero-truncated $\mathcal{N}^{+}(0,1)$ priors for the signal and error term variances.
[…]
3.3 Forecast combination based on estimates predictive ability
There are many reasons to model joint predictive ability. For example, it can be done purely for inference purposes, such as when we are interested in examining how the quality of predictions for different expert co-vary over a pooling space, or as a part of model evaluation. In this section, we focus on how to create a linear prediction pool (Geweke and Amisano,, 2011) based on these estimates.
A linear prediction pool can be made local by letting its weights depend on $\mathbf{z}$, so that the predictive density at a new point $\tilde{\mathbf{z}}$ in the pooling space is given by
$p\left(\tilde{y}\mid\tilde{\mathbf{z}}\right)=\sum_{k=1}^{K}w_{k}(\tilde{\mathbf{z}})p_{k}(\tilde{y}),\quad\sum_{k=1}^{K}w_{k}(\tilde{\mathbf{z}})=1,\quad w_{k}(\tilde{\mathbf{z}})\geq 0,$ (26)
where $w_{k}(\tilde{\mathbf{z}})$ is the weight of expert $k$ at $\tilde{\mathbf{z}}$. To bring the predictive ability of the expert into the pooling we let the weight $w_{k}(\tilde{\mathbf{z}})$ be a function of the posterior probability that expert $k$ has the highest predictive ability at $\tilde{\mathbf{z}}$. We denote this probability by $\psi_{k}(\tilde{\mathbf{z}})=p\big{(}\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}})\big{)}$, and refer to a forecast combinations based directly on $\psi_{k}(\tilde{\mathbf{z}})$ as $\operatorname{GP}(\nicefrac{{1}}{{3}})$ natural in tables and figures.
Often, and especially when the amount of data is small, these weights will tend to not discriminate strongly between experts and be outperformed by more aggressive weighing schemes such as the optimal linear pools of Geweke and Amisano, (2011). We therefore propose scaling $p\big{(}\eta_{k}(\tilde{\mathbf{z}})>\eta_{j\neq k}(\tilde{\mathbf{z}})$) by a constant $c$ and then feeding it through a softmax transformation
$w_{k}(\tilde{\mathbf{z}})=\frac{\exp\left(c\cdot\psi_{k}(\tilde{\mathbf{z}})\right)}{\sum_{j=1}^{K}\exp\left(c\cdot\psi_{j}(\tilde{\mathbf{z}})\right)}.$ (27)
@@ 3.3 Forecast combination based on estimates predictive ability
When the decision maker has no strong prior opinion of $c$, it can be dynamically selected at each time point by selecting the value that optimizes the historical predictive ability of the pool. We refer to this method as $\operatorname{GP}(\nicefrac{{1}}{{3}})$ dynamic in tables and figures. As a special case of interest, we will include $c=\infty$ under the name $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model selection, as it assigns weight 00 to all but one of the experts.
@@
Figure 9 shows how the optimal discrimination changes over time for the bike share data for the single-output method. At each time point, the historical performance of the pool for $0\leq c\leq 20$ is evaluated, and the next prediction is made with the value of $c$ corresponding to the historically most accurate pool. The optimal discrimination $c$ is volatile when little historical data is available, but then settles down around $c=5$ as a compromise between the extremes of equal weights and assigning all weight to the expert with highest probability of having the best predictive ability at $\tilde{\mathbf{z}}$.
The optimal discrimination factor for the multi-output model, illustrated in Fig. 10, settles around a lower value than in the single output version. This is due to the single-output methods failure to capture the strong positive correlation between the predictive ability of the experts, which naturally leads to a lower degree of discrimination. This is discussed further in Sect. 3.5.
Figure 9: Change in optimal discrimination factor over time, single output.
Figure 10: Change in optimal discrimination factor over time, multi-output.
All pooling methods described above will work regardless of whether the estimates of local predictive ability have been obtained with multiple single output GPs or a single multi-output GP, but taking account of the dependence will change the weights of the experts, sometimes drastically. This can be seen in Figs. 1113, which compare the single- and multi-output approaches for the three aggregation methods discussed above.
Figure 11: Cumulative log predictive scores based on the natural aggregation scheme for single and multi-output models.
Figure 12: Cumulative log predictive scores based on the model selection aggregation scheme for single and multi-output models.
Figure 13: Cumulative log predictive scores based on the dynamic aggregation scheme for single and multi-output models.
The multi-output model clearly outperforms its single output counterparts when using the natural weighting scheme. When using model selection or dynamic discrimination, however, the simpler single-output GP method performs slightly better, in both cases almost entirely due to a few observations at the end of the dataset.
3.4 Comparison with reference methods
@@
The sum of the log predictive scores over time for each method, together with the performance of the individual experts as well as a selection of reference methods, are shown in Table 2. The reference methods included are: the caliper method used in Oelrich et al., (2024), which is a nonparametric method that averages log scores of nearby observations in $\mathbf{z}$-space; the global optimal pool of Geweke and Amisano, (2011) which selects aggregation weights that optimize the historical performance of the pool in terms of log scores; the local optimal pool in Oelrich et al., (2024), which is an extension of the optimal pool of Geweke and Amisano, (2011) where the optimal weights are determined on a subset of the data in the pooling space closest to the current point. For details about these reference methods, see Oelrich et al., (2024).
While the natural multi-output GP outperforms its single-output counterpart considerably, this is reversed in the dynamic version. The single-output dynamic $\operatorname{GP}(\nicefrac{{1}}{{3}})$ performs the best out of the methods based on local predictive ability, but is soundly beaten by the local optimal pool of Oelrich et al., (2024).
Method Sum of log scores
SVBVAR 45.1
BREG 28.8
BART 43.6
Equal weights 98.7
Caliper 133.7
Global opt. 130.2
Local opt. 145.3
$\operatorname{GP}(\nicefrac{{1}}{{3}})$ natural 107.1
multi-$\operatorname{GP}(\nicefrac{{1}}{{3}})$ natural 115.9
$\operatorname{GP}(\nicefrac{{1}}{{3}})$ model sel. 126.6
multi-$\operatorname{GP}(\nicefrac{{1}}{{3}})$ model sel. 119.7
$\operatorname{GP}(\nicefrac{{1}}{{3}})$ dynamic 136.4
multi-$\operatorname{GP}(\nicefrac{{1}}{{3}})$ dynamic 131.7
Figure 14 compares the best performing of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ pooling schemes with the reference methods. The $\operatorname{GP}(\nicefrac{{1}}{{3}})$ initially struggles when there is little training data, but then quickly improves over time and at the end of the sample it only outperformed by the local optimal pool.
@@
3.5 Correlated experts and degree of discrimination
The multi-output model owes its superior performance when using natural weighting to a larger degree of discrimination, as the weak discrimination of the single-output method is in large part caused by a failure to capture correlation between the experts. A positive correlation that is unaccounted for will lead to weaker discrimination, and failure to take into account a negative correlation will lead to discrimination that is too strong. Typically, experts will be positively correlated and the degree of discrimination will be underestimated. This effect is illustrated in Fig. 15.
Figure 15: Gaussian contour plot illustrating the effect of correlation in predictive ability on the posterior probability that Expert 2 has greatest predictive ability at $\tilde{\mathbf{z}}$. In the top row, the marginal difference in posterior mean of predictive ability between the two experts is small ($\operatorname{E}(\eta_{2}(\tilde{\mathbf{z}}))-\operatorname{E}(\eta_{1}(\tilde{\mathbf{z}}))=0.5$), and in the bottom row the difference is substantial ($\operatorname{E}(\eta_{2}(\tilde{\mathbf{z}}))-\operatorname{E}(\eta_{1}(\tilde{\mathbf{z}}))=2$). The blue part of each contour plot marks the parts of the joint density where $\eta_{2}(\tilde{\mathbf{z}})>\eta_{1}(\tilde{\mathbf{z}})$.
Figure 15 displays contour plots of the posteriors for $\eta_{1}$ and $\eta_{2}$ for two experts in two setups. In the top row, the difference in posterior mean of predictive ability between the two experts is small (${\operatorname{E}(\eta_{2})-\operatorname{E}(\eta_{1})=0.5}$), and in the bottom row the difference is more substantial ($\operatorname{E}(\eta_{2})-\operatorname{E}(\eta_{1})=2$). In both setups we fix the marginal posterior variances of the two experts at $1$, and explore what happens when we let the correlation go from strong and negative $(\rho=-0.9)$ to non-existent $(\rho=0)$ to strong and positive $(\rho=0.9)$. As the correlation between the experts approaches one, the posterior probability of having the highest predictive ability will concentrate on the better model, no matter how small the difference in posterior means. Compared to uncorrelated experts, negative correlation increase the posterior probability that the expert with lower posterior mean has the higher predictive ability, something which becomes clearer as the magnitude of the difference in mean increases.
The effect illustrated in Fig. 15 explains why the multi-output GP outperforms the single output version when using natural weights. By capturing the correlation between experts, the multi-output GP leads to natural weights that discriminate more strongly between the experts.
@@ 4 Conclusions and further research
Gaussian processes have the drawback of being computationally heavy, and so we suggest the use of a power transformation of log scores which allows us to use a much faster marginalized GP to generate draws from the posterior of $\eta(\mathbf{z})$. Our proposed method is shown to work well for predicting bike share data (Fanaee-T and Gama,, 2014), where it outperforms all compared methods except the local version of the linear prediction pool of Geweke and Amisano, (2011) proposed in Oelrich et al., (2024).
Our method can be used to both model the marginal distribution of the predictive ability $\eta_{k}(\mathbf{z})$ for each expert, and to model the predictive ability of the experts jointly using a multi-output GP that makes it possible to capture correlation between the predictive abilities of the experts as well as correlation in the noise.
While the inference and prediction from the marginalized version is much faster compared to sampling also the latent variables in the GP($\nicefrac{{1}}{{3}}$) model, the step from running several single output GPs side by side to running a single multi-output GP increases the computational burden significantly, making large sample sizes time consuming. A topic for further study is therefore methods for speeding up the calculation using sparse linear algebra, methods for large-scale homoscedastic Gaussian GPs (Liu et al.,, 2020), or using alternative Bayesian inference approaches than Monte Carlo sampling, such as variational inference based on optimization (Blei et al.,, 2017).
Another potential avenue of future research is using alternative methods to pool predictions based on local predictive ability. In the empirical application we use three different methods, all of which have pooling weights based on $\psi_{k}$, the posterior probability that a particular expert has greater predictive ability than all other experts. Alternatively, one can form linear combinations directly on the $\eta_{1}(\mathbf{z}),\ldots,\eta_{K}(\mathbf{z})$ to obtain a posterior for the pooled local ability $\sum_{k=1}^{K}w_{k}\eta_{k}(\mathbf{z})$. Given a set of criteria for this distribution, such as maximizing the mean while keeping the variance below a certain threshold, a set of local weights can be obtained through numerical optimization.
@@
Allowing for time-varying weights has been shown to be beneficial in macroeconomic forecasting (Del Negro et al.,, 2014; Billio et al.,, 2013; McAlinn et al.,, 2020). While we have not explicitly discussed modeling the persistence of local predictive ability over time, it is straight forward to include some function of time in our vector of pooling variables via an appropriate kernel function.
Appendix
In this appendix we investigate the properties of our methods on simulated data. We are particularly interested in investigating how the computationally faster $\operatorname{GP}(\nicefrac{{1}}{{3}})$ model compares to the computationally costly $\operatorname{GP}(\chi_{1}^{2})$ model.
We generate data from the model
We further assume that our decision maker believes that the predictive ability of the expert varies over the variable $x_{2}$, but does not know the exact relationship between the ELDP $\eta$ and $x_{2}$. We will examine how well the decision maker will be able to estimate the true predictive ability, as summarized by the posterior distribution of $\eta$, using the two models described in Sect. 2.
The true distribution of the log scores in this setup is given by
[…]
Refer to caption Refer to caption
Figure 17: The left figure shows the difference in log mean integrated squared error (MISE) of the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ and $\operatorname{GP}(\chi^{2}_{1})$ model, taken over the distribution of $x_{2}$. The right figure shows the difference in mean integrated log score (MILS) of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ and $\operatorname{GP}(\chi^{2}_{1})$ model, taken over the distribution of $x_{2}$.
To compare the $\operatorname{GP}\left(\chi^{2}_{1}\right)$ and $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ models we simulate $1000$ data sets of $150$ observations each from the DGP in Eq. (28). For each simulation, we then generate draws from the posterior distribution of $\eta(x_{2})$ over a grid of $x_{2}$ values between $-3$ and $3$, and compare with the true ELPD as a function of $x_{2}$. Both GPs use a squared exponential kernel (Rasmussen and Williams,, 2006). We use $l\sim\operatorname{Inverse-Gamma}(5,5)$ priors for the length scales, and a $\alpha\sim\mathcal{N}^{+}(0,1)$ prior for the signal variances. The noise variance $\sigma_{n}^{2}$ of the $\operatorname{GP}(\nicefrac{{1}}{{3}})$ is given a $\mathcal{N}^{+}(0,1)$ prior as well.
Figure 16 show the posterior mean of $\eta(x_{2})$ together with a $95\%$ credible interval for three different data sets. The data sets have been selected by ranking the mean integrated squared error (MISE) of the ELPD estimates of each run—the MSE integrated over the distribution of $x_{2}$—and then selecting the $2.5$:th, $50$:th, and $97.5$:th percentile. Both methods perform well for $x_{2}\in(-2,2)$, but poorly in the tails. This is to be expected, as two factors conspire to make estimation in the tails difficult: the DGP variance of the log scores increases with the square of $x_{2}$, and since $x_{2}\sim\mathcal{N}(0,1)$ only roughly $5\%$ of the observations are in the tails.
Note that the uncertainty around the estimate increases quickly for both models when there are no, or few, previous predictions nearby. This will lead to the desirable behavior that when combining experts in regions where we have little to go on in terms of previous predictions, we will quickly tend towards an equal-weights solution; see Sect. 3.
The $\operatorname{GP}(\chi^{2}_{1})$ is prohibitively costly from a computational standpoint once the number of observations goes above a few hundred. That the more scalable $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model works so well compared to the $\operatorname{GP}(\chi^{2}_{1})$ model is therefore an important finding for practical applications. There is also a large collection of methods for reducing computational cost in GP regression with additive homoscedastic errors (Liu et al.,, 2020) that can be directly used to speed up inference in the $\operatorname{GP}\left(\nicefrac{{1}}{{3}}\right)$ model.